OpenStudy (anonymous):

Is the answer to this question true or false? Each side of the triangle shown has the same length. If the perimeter of the triangle is 49 inches, the length of each side is 16 1/3 inches.

OpenStudy (anonymous):

TRUE

OpenStudy (anonymous):

A triangle has 3 equal sides, so divide the perimeter by 3. Is it true or false?

OpenStudy (anonymous):

 x+x+x=3x=49 x=49/3=16 1/3

OpenStudy (anonymous):

Thanks